### v2

# From Trace to Line: LLM Agent for Real-World OSS Vulnerability Localization

Abstract
Large language models show promise for vulnerability discovery, yet prevailing methods inspect code in isolation, struggle with long contexts, and focus on coarse function or file level detections which offers limited actionable guidance to engineers who need precise line-level localization and targeted patches in real-world software development. We present T2L-Agent (Trace-to-Line Agent), a project-level, end-to-end framework that plans its own analysis and progressively narrows scope from modules to exact vulnerable lines. T2L-Agent couples multi-round feedback with an Agentic Trace Analyzer (ATA) that fuses runtime evidence such as crash points, stack traces, and coverage deltas with AST-based code chunking, enabling iterative refinement beyond single pass predictions and translating symptoms into actionable, line-level diagnoses. To benchmark line-level vulnerability discovery, we introduce T2L-ARVO, a diverse, expert-verified 50-case benchmark spanning five crash families and real-world projects. T2L-ARVO is specifically designed to support both coarse-grained detection and fine-grained localization, enabling rigorous evaluation of systems that aim to move beyond file-level predictions. On T2L-ARVO, T2L-Agent achieves up to 58.0% detection and 54.8% line-level localization, substantially outperforming baselines. Together, the framework and benchmark push LLM-based vulnerability detection from coarse identification toward deployable, robust, precision diagnostics that reduce noise and accelerate patching in open-source software workflows. Our framework and benchmark are publicly available as open source at https://github.com/haoranxi/T2LAgent.
1 Introduction
Software vulnerabilities now occur at unprecedented scale and cost. In 2023, more than 29,000 Common Vulnerabilities and Exposures (CVEs) were recorded cve_details_2024. In the first half of 2025, 1,732 data breaches were reported–an 11% increase indusface_2024. The economic toll is mounting: software supply-chain attacks are projected to cost the global economy 80.6 billion annually by 2026, up from 45.8 billion in 2023 darkreading_supply_chain; some estimates put total damages at \$9.5 trillion in 2024 liu_makridis_galinkin_2024. Critically, 14% of breaches in 2024 began with vulnerability exploitation—nearly triple the prior year. Despite advances in automated vulnerability detection, effectiveness in localization remains underexplored zhang2024empirical, leaving developers with coarse, file-level predictions while 32% of critical vulnerabilities remain unpatched for over 180 days. These vulnerabilities are exploding in volume and growing ever more severe, driving escalating security and economic risks worldwide.
Motivation. The rise of Large Language Models (LLMs) has accelerated AI-driven automation across software development and cybersecurity zhang_et_al_2025. From code assistants like github_copilot_2021 to automated security analysis and automation nunez2024autosafecoder. LLMs show strong aptitude for understanding complex codebases and flagging potential issues divakaran2024llms. This momentum naturally extends to vulnerability detection, spanning model-level approaches such as fine-tuning LLMs for classification—and agentic frameworks that harness LLM reasoning for automated analysis awesome_llm4cybersecurity_2024.
Yet today’s LLM-based detection methods face practical barriers. Most operate at the function level, asking whether a fragment is vulnerable rather than pinpointing where the flaw lies zhang2024empirical, 10.1145/3715758. These tasks are also defined purely on static code, ignoring binaries, runtime behavior, and project-level diffs, further limiting their usefulness for real remediation. Evaluations also rely heavily on lightweight, synthetic datasets that miss the complexity of production systems springer_vulnerability_analysis.
This research–practice gap is most visible in vulnerability localization. While studies show promising results on isolated functions or snippets, engineers must navigate large repositories with cross-module dependencies and need line-level localization to craft minimal, targeted patches precision current methods rarely deliver. Moreover, artificial benchmarks obscure the true difficulty of project-level detection, where context spans multiple files and demands system-wide reasoning.
Given the surge in vulnerability reports outlined earlier, the field urgently needs approaches that close this reality gap: leveraging LLM automation while tackling project-scale challenges practitioners face daily. Only by embracing these real-world constraints can we convert the promise of LLM-assisted security into tools that meaningfully lighten the load on development teams.
To address these gaps, we introduce T2L (Trace-to-Line), which reframes vulnerability detection as a two-tier problem (a) coarse-grained detection - flagging suspicious code chunks and (b) fine-grained localization - pinpointing exact vulnerable lines. This separation enables systematic evaluation of LLM capabilities from repository-scale reasoning to human-expert precision.
Contribution. This work makes three contributions: (1) New Task formulation. We define a new task for agentic AI, a runtime trace guided, project-level vulnerability detection, and formulate it as two structured subtasks: chunk-level detection and line-level localization. AST-based chunking adapts large codebases to LLM context while preserving semantics, enabling refinement from coarse predictions to exact lines, bridging research setups and real-world needs for LLMs. (2) T2L-ARVO benchmark. We present the first benchmark for agentic fine-grained localization, featuring 50 expert verified cases across five vulnerability types with balanced category distribution. T2L-ARVO enables realistic, project-scale evaluation of LLM-based systems. (3) T2L-Agent framework. We propose a multi-agent system with the following innovation: (i) a Trace Analyzer (Agentic Trace Analyzer) integrating the ensemble of tools such as static analysis, sanitizers, and runtime monitoring for observability as human debuggers; (ii) a Proposal module (Divergence Tracing) that iteratively forms and tests vulnerability hypotheses; and (iii) a two-stage pipeline (Detection Refinement) that first detects coarse chunks and then verifies exact lines with runtime feedback.
Overall, T2L-Agent provides a comprehensive baseline for this new vulnerability localization task. Unlike prior snippet-level detectors, it is designed to operate on fully reproducible real-world projects, using crashes, sanitizer reports, and project diffs as runtime evidence. On T2L-ARVO, it achieves up to 58% chunk-level detection and 54.8% exact line localization, showing that this runtime-evidence–guided setting is challenging yet tractable for LLM agents.
2 Background
Vulnerability Localization. Classical localization combines static and dynamic analyses with retrieval and graph methods. Static slicing selects statements that may affect a slicing criterion to form behavior-preserving slices for debugging and comprehension 10.5555/800078.802557. Dynamic slicing refines this by deriving input-specific slices from execution-time dependence graphs, yielding higher precision near crashes or variables agrawal1990dynamic. Information-retrieval approaches treat bug reports as queries, ranking files/methods by textual and historical signals to surface likely locations zhou2012should; 6693093; 10.1145/2597008.2597148. Code Property Graphs unify classic program-analysis concepts in a joint representation, enabling scalable pattern searches that uncovered 18 previously unknown vulnerabilities yamaguchi2014modeling.
AI for Cybersecurity. AI is widely applied to cybersecurity tasks with tool execution and feedback in runnable environments yang2023intercode. One prominent use is offensive security shao2024empirical, where agents reason about attack chains and interact with benchmark environments shao2024nyu; zhang2024cybench to simulate human CTF players udeshi2025d; shao2025craken; abramovichenigma. In parallel, pen-testing systems orchestrate modules or multi-agent collaboration to automate reconnaissance, vulnerability analysis, and exploitation against targets deng2024pentestgpt; shen2025pentestagent. Further work builds agent-based patching frameworks that localize vulnerabilities and synthesize fixes yu2025patchagent; xue2025pagent, and explores LLM-augmented red and blue teaming, where agents emulate adaptive attackers and assist defenders in threat hunting and incident investigation abuadbba2025promise; liu2025benchmarking.
LLM Agentic Systems. Beyond cybersecurity, large language models (LLMs) are increasingly applied to automate workflows across diverse domains shao2024survey, including AI for scientific research bran2023chemcrow; basit2025pennylang; jin2024prollm, software engineering yang2024swe, healthcare isern2016systematic, and hardware design wang2024llms. Many systems integrate long-term memory to retain intermediate results, environment states, and user preferences across sessions zhang2025survey. LLM as judge are often used to evaluate actions, select among tool outputs, and provide feedback for self-improvement  li2025generation. Some agents employ self-reflection and critique loops to iteratively refine their reasoning or code before execution renze2024self. Multi-agent and role-specialized architectures further enable collaboration among agents with distinct skills to solve complex, interleaved subtasks in parallel li2024survey.
Figure 1: T2L-Agent Framework overview.
3 Related Work
LLMs show strong potential on vulnerability localization tasks. Early learning-based detectors improve localization by training models to classify whether a unit is vulnerable. LLMAOyang2024large fine-tunes LLMs on small, manually curated buggy programs, while BAPstein2025s learns state-of-the-art vulnerability localization without any direct localization labels, outperforming traditional baseline over eight benchmarks. However, subsequent analyses revealed data issues in widely used datasets such as Big-Vul and Devignzhou2019devign, making people doubt about performance numbers and highlighting the need for realistic evaluation settings croft2023data.
Study
Line Lv.
Mult. Ag.
Runtime
Iterative
LLMAO yang2024large
BAP stein2025s
GenLoc asad2025leveraging
AgentFL qin2024agentfl
CoSIL jiang2025cosil
AutoFL kang2024quantitative
LineVul 9796256
LOVA li2024attention
MatsVD 10.1145/3671016.3674807
xLoc 10.1145/3660804
LLM4FL rafi2024multi
MemFL yeo2025improving
T2L (ours)
Figure 2: Related Works
Many prior works frame vulnerability localization at the file or function level, but this coarse granularity often fails to offers limited guidance for developers. GenLoc asad2025leveraging identifies potentially vulnerable files from bug reports and iteratively analyzes them using code exploration tools. AgentFL qin2024agentfl applies a multi-agent framework for function-level localization, modeling the task as a three-step pipeline with specialized agents and tools. CoSIL jiang2025cosil narrows the function-level search space using module call graphs and iteratively traverses them for relevant context. Similarly, AutoFL kang2024quantitative prompts LLMs to localize method-level vulnerabilities via function-call navigation, showing that multi-step reasoning mitigates context limits.
Recent studies have shifted toward line or statement level localization. LineVul 9796256 uses a Transformer-based classifier for line-level prediction, while LOVA li2024attention introduces a self-attention framework to score and highlight vulnerable lines. MatsVD 10.1145/3671016.3674807 enhances statement level localization with dependency-aware attention, and xLoc 10.1145/3660804 learns multilingual, task-specific knowledge for bug detection and localization. Building on these efforts, LLM4FL rafi2024multi proposes a multi-agent framework leveraging graph-based retrieval and navigation to reason about failure causes. MemFL yeo2025improving introduces external memory to incorporate project-specific knowledge, improving localization in complex, repository-scale systems. CThey push localization to line level and use more multi-agent coordination. However, most still rely on limited runtime evidence, single-pass predictions, or benchmarks that lack realistic project settings.
T2L formalizes a new task of runtime-trace-guided vulnerability localization at the project level. Prior learning-based approaches such as BAP, LLMAO, and LineVul rely on static code analysis, training models to classify whether a given code snippet is vulnerable based solely on syntactic and semantic patterns. These methods typically operate on isolated single files and produce coarse-grained predictions at the function or file level without leveraging dynamic execution information. In contrast, our task targets multi-file codebases and incorporates runtime evidence including reproducible crash logs, stack traces, and sanitizer reports to pinpoint exact vulnerable lines. This formulation more closely reflects real-world vulnerability triage workflows where developers rely on both static code inspection and dynamic program behavior to diagnose security flaws.
4 Methods
4.1 T2L Agent
T2L-Agent uses a hierarchical planner–executor design udeshi2025d that splits localization into evidence gathering, hypothesis generation, and iterative refinement.Unlike single-pass analyzers, it uses a human-like process: gather runtime signals, link them to code, and narrow the search space. Details of the tool list that T2L originally supports are provided in  A.5.
Figure 3: ATA Components
Evidence Tracing T2L Planner coordinates repository analysis and runtime evidence capture in a single, structured pipeline. First, code-structure analysis partitions the codebase into function-aligned, semantically coherent units that preserve syntactic relationships while fitting LLM context windows; known patch locations are also indexed to establish evaluation baselines. Next, runtime evidence becomes the cornerstone: Sanitizer records memory-violation patterns, allocation traces, and stack frames, while interactive debugging (GDB/LLDB) provides symbolic context via backtraces and variable-state snapshots at crash points. This dual-layer design yields comprehensive, observable crash logs rather than speculative static signals. We collect multiple types of runtime and static evidence, including crash logs, sanitizer reports, stack traces, and static analyzer results, all evidence is merged into a single unified textual block and fed to the LLM at once. This design ensures that the LLM forms its reasoning based on the global evidence context, rather than aggregating per-evidence predictions to avoids issues of cross-signal inconsistency. The integrated workflow is a key T2L innovation; we detail it in Sec. 3.
Hypothesis Generation A “hypothesis” an LLM-generated candidate vulnerability produced from the unified evidence block of an existing vulnerable project in T2L system. The LLM proposes a list of file:line candidate pairs ranked by confidence. Each hypothesis is independently checked against the ground-truth patch diff, and candidates do not affect one another during evaluation.
Two-stage Refinement To ensure accurate detection, T2L uses a two-stage refinement process. It begins with coarse-grained filtering: the system derives broad candidate regions from crash logs, using LLM code comprehension to link symptoms to likely causes and produce a ranked list of candidates with confidence and rationale. It then identifies code fragments the LLM flagged as relevant in this first pass–such as missing bounds checks, uninitialized variables, or improper memory management. Next comes fine-grained refinement: these extracted fragments are added to the evidence for a second LLM pass, which updates the candidate set. Any newly generated locations are merged into the existing candidates. This refinement loop continues until the LLM produces no additional candidates or the cost budget is reached.
Feedback Control Each cycle outputs a brief task summary with success indicators and confidence. The Planner adapts the next step—continue refining or stop, preventing premature termination and avoiding over-analysis. The loop improves precision while controlling compute cost.
Agentic Trace Analyzer (ATA) The Agentic Trace Analyzer (ATA) is the core module of T2L-Agent. Its purpose is to connect runtime behavior with relevant source code regions, addressing a gap in prior work that relies mainly on static signals. ATA constructs a unified “evidence block” by running targets in reproducible Docker, partitioning repositories into semantically meaningful chunks using tree-sitter, and instrumenting executions with analysis toolkit such as Sanitizers, Debuggers, and Static analyzers to collect stack traces, memory errors, and other runtime cues. The output is a unified “evidence block” which is then passed to T2L-Agent’s reasoning modules.
We use hierarchical refinement, starting from crash logs and narrowing to specific code. The agent cross-validates runtime signals with static features and ranks candidates by LLM confidence. Like human debugging, T2L-Agent seeds candidates via static–dynamic correlation and iteratively refines them with source–feedback loops. ATA brings fewer single-shot failures, improved compute efficiency, and behavior-anchored decisions that enable precise, line-level localization even in large, tightly coupled codebases. With ATA disabled, the LLM localizes on its own, as shown in Figure 4(c). Without ATA, the LLM fails to localize vulnerabilities compared with (a) and (b).
(a) Detection Refinement
(b) Divergence Tracing
(c) Agentic Trace Analyzer
Figure 4: Partial T2L-Agent logs to show the how the three proposed technique on T2L-Agent work and help the task: Detection Refinement, Divergence Tracing and Agentic Trace Analyzer.
4.2 T2L Features
Divergence Tracing. Recognizing that complex vulnerabilities often involves multiple files and functions, T2L-Agent also uses divergence tracing to explore multiple hypotheses in parallel from the same crash log. This feature is inspired by how modern LLM interfaces offer multiple response variations for users, using the LLM’s variability to ensure comprehensive coverage of potential vulnerability locations. Rather than committing to a single chain of thought, it expands several in parallel and returns a ranked list of candidate sites across the search space. This surfaces correct localizations that were not top ranked initially and is especially helpful for bugs that span multiple modules. From Figure 4 (b), we can observe that the divergence tracing generates more localization candidates and matched more vulnerable lines in this round.
Detection Refinement. The detection refinement begins based on crash logs and initial localization candidates. Each LLM-generated candidate gets a confidence label from very unlikely to very likely. We use these labels to sort and rank the candidates during the coarse-grained filtering. T2L agent selects code slices based on crash logs, stack trace information, and vulnerability patterns instead of exhaustively examining the full code snippet. In fine-grained refinement, we extract these slices and feed them back to the LLM along with the unified evidence, providing richer context surrounding the initially flagged locations.
The refinement process operates iteratively to help the agent correct early mistakes and discover vulnerabilities that are not immediately obvious from traces, particularly for complex vulnerabilities involving memory corruption where the crash point may be far away from the actual vulnerability. As Figure 4 (a) shows, based on the source code LLM interested, the refinement process successfully locates new lines that were not found during first step with only the runtime evidence.
4.3 T2L-ARVO Benchmark
The original ARVO dataset contains over 4,993 reproducible vulnerabilities across 250+ C/C++ projects, but its human-oriented design and build-centric structure do not directly support evaluating agentic vulnerability-localization systems. To bridge this gap, we developed T2L-ARVO, a 50-case benchmark derived from ARVO for assessing LLM agents in evidence-guided vulnerability localization, rather than human bug reproduction.T2L-ARVO maintains a moderate, calibrated difficulty while keeping the original crash-type diversity. Its construction employs a dual validation pipeline combining manual expert review with LLM-assisted checks to ensure each case is reproducible and appropriately challenging for automated agents. This yields a realistic, balanced benchmark for rigorous end-to-end evaluation of LLM-based vulnerability localization. In T2L-ARVO, each of the cases has multiple ground-truth bug lines. The average accuracy overall is denoted as $\textit{Accuracy}=\textit{correct localized vulnerable lines}/\textit{total vulnerable lines}$.
T2L-ARVO Composition We analyzed 4,9934{,}993 ARVO instances and grouped them by underlying failure mechanism: Buffer Overflows $49.9\%$ (n=2,490)(n=2{,}490), Uninitialized Access & Unknown States $35.4\%$ (n=1,768)(n=1{,}768), Memory Lifecycle Errors $11.5\%$ $(n=573)$, Type Safety & Parameter Validation $2.9\%$ $(n=147)$, and System & Runtime Errors $0.3\%$ $(n=15)$. Each family subsumes concrete subtypes (e.g., heap-buffer-overflow, use-of-uninitialized-value, heap-use-after-free, bad-cast). T2L-ARVO deliberately mirrors this distribution to avoid bias toward any single failure mode.
The final benchmark comprises 50 vulnerabilities samples, evenly sampled across five crash families (10 each) for broad yet controlled difficulty. Each family includes representative subtypes (e.g., heap-buffer-overflow, use-of-uninitialized-value, heap-use-after-free, bad-cast), covering single-file defects and cross-module interactions to prevent overflow bias and exercise diverse failure modes observed in real repositories.
Figure 5: Crash types in T2L-ARVO Bench.
Crash Family Brief Description
Buffer Overflow Violations of memory bounds (heap/stack)
Uninitialized Access Reads from undefined or indeterminate state
Memory Lifecycle Use-after-free / double-free / lifetime bugs
Type Safety Bad casts, invalid args, contract violations
System Runtime Environment and runtime interaction faults
Verification Process Our verification follows two main considerations. First, ARVO crash families contain multiple subtypes. To reduce redundancy and improve representativeness, we select only a few typical subtypes from each family (see A.1), preserving diversity without overloading the benchmark with similar vulnerabilities. Second, experts evaluate each sample’s complexity to keep T2L-ARVO at a moderate, calibrated difficulty. We score candidates using diff-based structural metrics (e.g., files changed, architectural spread, directory depth) and semantic factors (e.g., cross-module coupling, interface changes), ensuring that T2L-ARVO remains balanced in both coverage and difficulty.
5 Experiment Setup
Metrics. We report two complementary scores for project-level OSS vulnerability studies. Detection asks whether the agent flags a vulnerability within the correct module/chunk and materially shrinks the search space. Localization requires exact line matches to ground-truth patches. Together, they separate “finding the neighborhood” from “pinpointing the line”, mirroring real debugging.
Data Preparation. We evaluate on the full T2L-ARVO set: 50 verified, structured challenges derived from ARVO, spanning diverse domains (e.g., imaging, networking) and balanced complexity so results reflect production patterns rather than a single project or bug family. Because ARVO lacks detection-ready chunking, T2L-ARVO adds AST-based segmentation: projects are partitioned into semantically meaningful units for scoring coarse detection, while exact line matches assess fine localization—a single framework for both levels.
Model Selection. We assess a set of state-of-the-art language models—both open-source and commercial to probe generality and robustness of T2L-Agent across architectures and scales, including open models such as Qwen3 Next, Qwen3 235B, DeepSeek 3.1, LLaMA 4 and commercial models like Claude4 Sonnet, GPT-5, GPT-4.1, GPT-4o-mini, Gemini 2.5 Pro, Gemini 2.5 Flash with a maximum budget \$1.0. We use API keys from commercial model’s official providers and Together.ai’s inference service for open source models.
Implementation. We build T2L-Agent from scratch without LangChain, DSPy, and LlamaIndex to keep the core lightweight, retain fine-grained control over reasoning and tools, and maximize extensibility. The Agentic Trace Analyzer compiles targets with ASAN and collects crashes, stack traces, and allocation metadata to yield actionable traces for narrowing. Following ARVO’s layout, we maintain registry-backed, per-project environments and provide both vulnerable and patched revisions in containers. Our harness runs dockerized T2L-Agent’s that interface with T2L-ARVO via the Docker SDK for Python, orchestrating build–run–reproduce cycles from within the agent loop, ensuring consistent conditions, deterministic reproduction, and auditable measurement.
6 Evaluation
6.1 Baseline Benchmarking
Table 1: Base T2L-Agent Localization and Detection Rate Performance Across Different Models.
% Avg. Buffer Initialize Memory Parameter Runtime
Det Loc Det Loc Det Loc Det Loc Det Loc Det Loc
GPT-5 44.3 41.7 57.5 53.8 35.6 35.5 60.8 55.9 36.5 39.4 11.2 10.0
GPT-4.1 48.0 38.5 60.8 36.5 50.6 46.4 60.8 46.1 26.5 29.9 21.2 20.5
GPT-4o-mini 44.3 22.6 60.8 20.2 48.1 12.5 55.8 24.7 28.7 26.7 11.2 10.0
Claude 4 Sonnet 45.9 30.5 57.5 50.6 60.8 36.1 1.3 31.6 37.8 46.1 24.8 0.5
Gemini2.5 Pro 17.4 10.5 25.0 5.6 25.0 20.0 11.3 10.8 5.4 11.9 10.0 10.5
Qwen3 235B 25.9 9.2 25.8 6.5 23.1 1.7 40.8 16.7 28.7 17.3 7.9 0.0
Qwen3 Next 80B 37.4 5.9 54.2 3.7 33.1 0.4 55.8 14.5 29.1 6.8 1.2 0.0
We evaluate T2L-Agent on T2L-ARVO end-to-end and report Detection Rate and Localization Rate. Table 1 covers five models—GPT-5, GPT-4.1, GPT-4o-mini, Qwen 3 Next 80B, and Qwen3 235B, running under identical per-case budgets, environments, and AST-based chunking. Overall, detection is higher than localization by design. Under this setting, GPT-5 leads localization at 41.7%, followed by GPT-4.1 at 38.5%; GPT-4o-mini lands at 22.6%, and open-source models trail Qwen 3 235B 9.2% and Qwen 3 Next 80B 5.9%. For detection, GPT-4.1 is highest at 48.0%, with GPT-5 and GPT-4o-mini both at 44.3%; Qwen3 reaches 37.4% and Qwen3 235B 25.9%. While Gemini2.5 Pro shows limited effectiveness with 17.4% detection and 10.5% localization rates.
Family-wise patterns are consistent across metrics. Buffer and Memory are easier due to concrete runtime cues: for localization, GPT-5 reaches 53.8% and 55.9%, and most models cluster near the mid-50s for detection. Initialize sits mid-range and benefits from multi-step reasoning (e.g., GPT-5 35.5% vs. GPT-4.1 46.4% in localization). Parameter is often solvable from interface/call-site context (39.4% for GPT-5; 29.9% for GPT-4.1). Runtime remains uniformly hardest: detection hovers around 11.1–21.2% even for top configurations, and the best localization we observe is 20.5% on GPT-4.1, reflecting sparse, unstable traces.
Taken together, equal budgets surface clear profiles. GPT-5 (low-think) converts evidence into the strongest line-level localization, while GPT-4.1 extracts slightly more coarse-grained signal at the chunk level. GPT-4o-mini’s mix—competitive detection (44.3%) but weak localization (22.6%), suggests higher recall with looser ranking that does not always translate to precise line hits. Open-source models lag on both metrics under the same constraints, indicating gaps in code understanding and tool use rather than simple parameter tuning. Overall, improvements track the availability of concrete runtime evidence, and structured, tool-grounded reasoning appears more impactful than generation settings for end-to-end vulnerability localization.
6.2 Discussion 1: Feature-wise Evaluation
Agentic Trace Analyzer. This table  2 demonstrates the critical effectiveness of our proposed Agentic Trace Analyzer (ATA) through ablation experiments. Without ATA, GPT-5 and Claude 4 Sonnet show 0.0% detection and localization across all families. This performance breakdown validates that our ATA component successfully bridges the gap between crash symptoms and vulnerability locations, addresses the fundamental challenge of vulnerability localization in complex codebases.
Detection Refinement. Compared with Tab  1, Tab. 2 shows broad, across-the-board gains after enabling refinement. Strong proprietary models improve steadily, while open-source models jump the most—Qwen3 235B’s localization rises by roughly sevenfold. Improvements vary by crash family: Initialize bugs benefit most (they demand multi-step reasoning), whereas Buffer and Memory see smaller lifts because concrete runtime evidence already anchors the search. Runtime cases remain hard—when traces are sparse, refinement offers limited benefit. Net effect: higher recall and more precise line-level hits with minimal tuning. Several additional models show promising performance. Deepseek V3.1 achieves the highest overall results with 53.9% detection and 53.4% localization rate. LLaMa 4 demonstrates balanced capabilities on both metrics, and Gemini 2.5 Flash shows variable performance across crash families.
Table 2: Localization and Detection Rate Performance of T2L-Agent with Refinement and Divergence Tracing Across Different Models.
% Avg. Buffer Initialize Memory Parameter Runtime
Det Loc $\DeltaDet Δ\DeltaLoc$ Det Loc Det Loc Det Loc Det Loc Det Loc
w/ Detection Refinement
[…]
Gemini 2.5 Flash 22.5 18.4 34.2 0.6 40.8 25.7 7.9 27.6 0.4 33.4 24.6 0.6
Llama4 28.3 28.1 30.8 25.6 35.8 26.1 0.0 27.9 24.1 32.0 30.2 0.0
Deepseek V3.1 53.9 53.4 60.8 55.6 62.5 47.5 16.3 58.1 55.0 60.2 47.6 19.4
w/ Divergence Tracing
[…]
w/o Agentic Trace Analyzer
[…]
Divergence Tracing. Tab. 2 shows divergence tracing yields the strongest gains. All models improve on both metrics: GPT-5 is up 13.7% in detection and 10.3 in localization; GPT-4.1 gains 4.0% and 11.4% respectively. Qwen3 Next see the largest jumps with adding 13.8% in detection and 48.9% in localization, while Qwen3 235B adds 16.8% and 32.9%. These consistent lifts across architectures highlight divergence tracing as a core algorithmic upgrade for vulnerability localization.
6.3 Discussion 2: Parameter Tuning
Thinking Budget. As shown in Tab.  3, more thinking didn’t help. On GPT-5, the Medium budget outperforms High with 50.9% detection vs 41.3%, and 41.6% localization vs 36.1%. The Low setting trails Medium by a few points yet often matches or even exceeds High on key metrics, while sharply reducing compute and latency. This pattern suggests diminishing returns—and decision drag—at very high budgets: the model over-explores, delays commitment, and accumulates tool-use errors. In practice, Medium strikes the best accuracy–cost balance for vulnerability localization; Low is a strong option when throughput and responsiveness matter most.
Table 3: Localization and Detection Rate Performance of T2L-Agent for Temperature Tuning and Chain of Thought Across Different Models.
% Avg. Buffer Initialize Memory Parameter Runtime
Config Det Loc Det Loc Det Loc Det Loc Det Loc Det Loc
Temperature
GPT-4.1 0.2 51.0 43.6 57.5 44.4 55.6 51.1 55.8 45.7 43.2 35.4 21.2 20.2
GPT-4.1 0.6 50.8 43.5 60.8 52.0 53.1 47.8 55.8 44.8 39.4 42.7 17.9 10.1
Claude 4 Sonnet 0.2 46.5 43.0 54.2 56.1 55.6 43.0 61.7 49.0 36.5 39.7 11.2 10.9
Claude 4 Sonnet 0.6 47.3 44.9 54.2 56.1 50.6 48.0 60.8 53.6 36.5 39.6 11.2 10.9
Reasoning Effect
GPT-5 High 41.3 36.1 54.2 47.8 40.6 32.4 55.8 39.3 32.8 36.7 10.0 10.0
GPT-5 Medium 50.9 41.6 60.8 55.6 60.8 42.8 11.3 46.1 45.8 40.4 47.4 10.5
Temperature. Temperature changes barely matter from Tab.  3. On GPT-4.1, detection is 51.0% at 0.2 and 50.8% at 0.6, with localization 43.6% vs. 43.5%. Claude 4 Sonnet shows the same pattern: 46.5% vs. 47.3% detection and 43.0% vs. 44.9% localization. Performance is stable in 0.2-0.6 range. The exception is Initialize bugs, which are more temperature sensitive than Buffer and Memory cases that lean on concrete runtime evidence. Overall, precise localization benefits more from structured, tool-grounded reasoning than from extra sampling, making parameter choices simple.
6.4 Discussion 3: Case Study
Figure 6 illustrates a full T2L-Agent workflow on a real case from T2L-ARVO, showcasing the iterative planner–executor architecture in action. The process starts with the Planner orchestrating code chunking (chunk_case, 5441 chunks) and diff indexing (diff_index), then running sanitized execution (run_san) to collect crash logs before delegating reasoning to the Executor.
Figure 6: T2L-Agent pipeline visualization.
The Executor analyzes traces via llm_analyze, extracts ranked candidates (extract_json), and iteratively evaluates them against ground truth using compare_llm_metrics, combining static patterns and dynamic signals. This multi-round refinement achieves perfect localization (detection: 1.0, localization: 1.0), demonstrating T2L-Agent’s ability to convert crash symptoms into precise diagnostics. Each panel visualizes thought (function call) and observation (result) with hand-drawn borders and no hallucinated text. Together, they highlight data flow, role separation, and metric-driven validation across planner and executor components.
7 Limitation and Future Work
Our work has three key limitations. First, the T2L-ARVO benchmark includes only 50 manually verified cases. While it offers broad vulnerability coverage and balanced categories, the limited sample size constrains evaluation due to the human verification efforts. Also, ARVO’s dataset structure is designed for human developers, which needs more fine-grained metadata that could benefit LLM-based localization. Second, although T2L-Agent improves localization accuracy from 0% to 54.8% through three innovations, cost efficiency remains a concern. The agent operates effectively under a \$1.0 budget via task-aware planning and early stopping, but large-scale deployment across thousands of vulnerabilities would demand significant optimization. Third, higher model thinking budgets fail to boost localization performance, indicating that increased compute alone is insufficient. This points to a need for smarter ways to exploit model reasoning. Future work should explore more efficient architectures. Such as model cascading to coordinate cheaper and stronger models, and specialized multi-agent systems where roles are tailored to tools like our Agentic Trace Analyzer. These strategies may retain quality while scaling to production workloads.
8 Conclusion
T2L addresses a key gap between LLM-based vulnerability localization and real-world practice. We contribute three advances that move from coarse identification to precise diagnostics. First, we propose a new formulation for LLM-based vulnerability detection: chunk-wise detection and line-level localization, enabling structured and fine-grained evaluation. Second, T2L-ARVO introduces a benchmark for agentic line-level localization, with 50 expert-verified cases across diverse vulnerabilities. Third, T2L-Agent improves performance via our Agentic Trace Analyzer, which fuses runtime and static signals, as well as Divergence Tracing and Detection Refinement in a feedback-driven workflow. T2L-Agent achieves 44–58% detection and 38–54.8% localization, marking a step toward deployable systems for real-world code security.
Ethics Statement
This work builds upon the ARVO dataset for vulnerability analysis. Our T2L-ARVO benchmark is constructed using its full-version data, and we have properly cited the original ARVO project to acknowledge its contribution and comply with copyright and attribution standards. No additional data collection, user studies, or ethically sensitive procedures involved in the construction or evaluation of T2L-Agent. All experiments were conducted on Linux servers with only open source dependencies, and our system does not involve any privacy-sensitive data, bias-sensitive decision-making, or potentially harmful applications. We also note that large language models were used solely for light editing and polishing of manuscript, with no involvement in system design, code generation, or experimental results. Their use was limited to improving readability and presentation clarity.
Reproducibility Statement
To ensure reproducibility, we will publicly release the full T2L-Agent framework upon publish of this paper, including all module codebase, evaluation scripts, and benchmark data used in this paper. Our implementation does not rely on any proprietary components. Detailed descriptions of our methodology are provided in the main text (Sections 4.1, 4.3) and supported by step-by-step examples in the appendix. All experimental configurations, including model versions, prompting strategies, and budget constraints, will be documented and made available upon publication to enable full replication of our results.
Appendix A Appendix
A.1 ARVO Crash Type
To understand where our vulnerability localization agent should focus, we analyze the crash type distribution in ARVO dataset. ARVO’s crash type distribution is dominated by classic memory corruption: Buffer Overflow accounts for 49.9% of crashes, led by heap-buffer-overflow (36.1% overall) and followed by stack-buffer-overflow (6.2%), index-out-of-bounds (3.3%), and global-buffer-overflow (3.2%), with underflows/containers in the long tail (around 1% each). Uninitialized Access & Unknown States is the second largest family at 35.4%, primarily use-of-uninitialized-value (20.3%), then UNKNOWN READ/WRITE (around 11.8% combined). Memory Lifecycle Errors contribute 11.5%, dominated by heap-use-after-free (7.8% overall) plus double-free, use-after-poison, and invalid frees. Type Safety & Parameter Validation is smaller (2.9%)—notably bad-cast (1.3%) and negative-size-param (0.8%). System & Runtime Errors are rare (0.3%). Overall, around 85% of ARVO crashes fall into Buffer Overflow or Uninitialized/Unknown categories.
Table 4: Crash Families and Subtypes Analysis
Family Subtype Count % within family % of total
Buffer Overflow Vulnerabilities
Total 2490 49.9%
Heap-buffer-overflow 1802 72.4% 36.1%
Stack-buffer-overflow 308 12.4% 6.2%
Index-out-of-bounds 165 6.6% 3.3%
Global-buffer-overflow 160 6.4% 3.2%
Container-overflow 33 1.3% 0.7%
Stack-buffer-underflow 13 0.5% 0.3%
Dynamic-stack-buffer-overflow 9 0.4% 0.2%
Uninitialized Access & Unknown States
Total 1768 35.4%
Use-of-uninitialized-value 1015 57.4% 20.3%
UNKNOWN READ 462 26.1% 9.3%
Segv on unknown address 134 7.6% 2.7%
UNKNOWN WRITE 123 7.0% 2.5%
Null-dereference READ 25 1.4% 0.5%
UNKNOWN 8 0.5% 0.2%
Unknown-crash 1 0.1% 0.0%
Memory Lifecycle Errors
Total 573 11.5%
Heap-use-after-free 389 67.9% 7.8%
Heap-double-free 63 11.0% 1.3%
Use-after-poison 48 8.4% 1.0%
Invalid-free 29 5.1% 0.6%
Stack-use-after-return 26 4.5% 0.5%
Stack-use-after-scope 13 2.3% 0.3%
Bad-free 5 0.9% 0.1%
Type Safety & Parameter Validation
Total 147 2.9%
Bad-cast 65 44.2% 1.3%
Negative-size-param 42 28.6% 0.8%
Memcpy-param-overlap 20 13.6% 0.4%
Object-size 9 6.1% 0.2%
Incorrect-function-pointer-type 6 4.1% 0.1%
Non-positive-vla-bound-value 3 2.0% 0.1%
Strcpy-param-overlap 1 0.7% 0.0%
Strncpy-param-overlap 1 0.7% 0.0%
System & Runtime Errors
Total 15 0.3%
Check failed 6 40.0% 0.1%
Unknown signal 6 40.0% 0.1%
Bad parameters to –sanitizer-annotate-contiguous-container 2 13.3% 0.0%
Nested bug in the same thread, aborting. 1 6.7% 0.0%
Table 4: Crash Families and Subtypes Analysis (continued)
A.2 ARVO Dataset Profiling
We profiled the full ARVO corpus with all 4,993 vulnerabilities across 288 projects to guide T2L-ARVO’s design and document its coverage. The analysis maps distributional patterns, project traits, and crash-type frequencies, and clarifies how our 50-case subset aligns with the broader ARVO ecosystem. These profiles confirm that T2L-ARVO is representative across crash families, project complexity, and severity, providing a transparent baseline for extensions and alternative benchmarks. We include compact visualizations of these profiles to convey key ARVO factors—such as crash families, project complexity, and severity at a glance.
Figure 7: Distribution of vulnerability counts across the 288 ARVO projects. Most projects have under 25 vulnerabilities, with a long-tail of highly vulnerable ones.
Figure 8: Breakdown of vulnerability severities in ARVO. Over 70% are medium severity, while high severity cases account for the remaining 29%.
Figure 9: Analysis of fuzzing tools used in ARVO. libFuzzer dominates at 74.7%, followed by AFL and honggfuzz.
Figure 10: Visualization of the 30 most vulnerable projects in ARVO. ImageMagick leads with 368 vulnerabilities, with others showing diverse security footprints.
Figure 11: Project grouping by vulnerability count. Nearly half the projects have only 2–5 vulnerabilities, with very few exceeding 100.
A.3 Model Failure Analysis
We conducted a targeted failure analysis on several baseline models to map out common failure modes as shown in Tab.  12. Claude 4 Sonnet and Gemini 2.5 Pro hit the budget ceiling in 81.6% and 85.7% of runs respectively, indicating efficient resource utilization. GPT-5 reaches 61.2% with execution errors (28.6%), while Qwen3 235B struggles with basic data operations (59.2%). Open-source baselines stall early: Qwen 3 Next fails to surface actionable candidates in 44.9% of trials. Execution errors remain common across older models (20–30%), showing that tool use often breaks even when a plan exists. Net-net, while newer models show improved resource management, legacy models skew toward either incomplete exploration or difficulty navigating real-world code.
Figure 12: Model failure type distribution across five models on T2L-ARVO. GPT-5 commonly fails due to budget limits, while Qwen3 models often fail to generate actionable candidates.
A.4 T2L-ARVO Challenge List
We provide the comprehensive list of T2L-ARVO benchmark we verified and collected in this work along with the key meta information for each challenge.
Table 5: Bug Analysis Results by Category
Id Fuzzer Sanitizer Project Crash Type Severity
System & Runtime Errors
16737 libfuzzer ubsan graphicsmagick Unknown signal -
7966 libfuzzer ubsan graphicsmagick Unknown signal -
7654 libfuzzer ubsan graphicsmagick Unknown signal -
7639 libfuzzer ubsan graphicsmagick Unknown signal -
59193 libfuzzer msan faad2 Check failed -
49915 libfuzzer msan ndpi Check failed -
48780 libfuzzer msan libvpx Check failed -
7361 afl asan ots Bad parameters to sanitizer -
32939 libfuzzer asan rdkit Bad parameters to sanitizer -
Buffer Overflow Vulnerabilities
16614 libfuzzer asan opensc Heap-buffer-overflow Med
13956 libfuzzer asan yara Heap-buffer-overflow Med
16615 libfuzzer asan opensc Heap-buffer-overflow Med
20856 libfuzzer asan ndpi Heap-buffer-overflow Med
17330 libfuzzer asan openthread Stack-buffer-overflow High
42454 libfuzzer asan ghostpdl Stack-buffer-overflow High
17297 libfuzzer asan openthread Stack-buffer-overflow Med
18562 libfuzzer asan lwan Global-buffer-overflow -
30507 honggfuzz asan serenity Global-buffer-overflow -
18231 afl asan binutils-gdb Global-buffer-overflow -
Uninitialized Access & Unknown States
49493 libfuzzer asan mruby Segv on unknown address -
24290 honggfuzz asan libvips Segv on unknown address -
57037 libfuzzer asan mruby Segv on unknown address -
23778 libfuzzer msan binutils-gdb Use-of-uninitialized-value Med
20112 libfuzzer msan open62541 Use-of-uninitialized-value Med
47855 libfuzzer msan harfbuzz Use-of-uninitialized-value Med
16857 libfuzzer msan matio Use-of-uninitialized-value Med
43989 libfuzzer asan ghostpdl Null-dereference -
2623 libfuzzer asan h2o Null-dereference -
45320 libfuzzer asan ghostpdl Null-dereference -
Memory Lifecycle Errors
42503 libfuzzer asan php-src Heap-use-after-free High
38878 libfuzzer asan harfbuzz Heap-use-after-free High
14245 afl asan karchive Heap-use-after-free High
19723 libfuzzer asan leptonica Heap-use-after-free Med
33750 honggfuzz asan fluent-bit Heap-double-free High
34116 honggfuzz asan fluent-bit Heap-double-free High
20785 libfuzzer asan llvm-project Use-after-poison High
3505 afl asan librawspeed Use-after-poison High
51687 afl asan mongoose Use-after-poison High
31705 afl asan c-blosc2 Invalid-free -
Type Safety & Parameter Validation
2798 libfuzzer ubsan gdal Bad-cast High
29267 libfuzzer ubsan serenity Bad-cast High
33150 libfuzzer ubsan libredwg Object-size Med
20217 libfuzzer ubsan arrow Object-size Med
12679 afl asan openthread Memcpy-param-overlap Med
23547 honggfuzz asan php-src Memcpy-param-overlap Med
25357 libfuzzer asan libsndfile Negative-size-param -
60605 libfuzzer asan ndpi Negative-size-param Med
2692 libfuzzer ubsan boringssl Incorrect-function-pointer-type Med
50623 libfuzzer ubsan serenity Non-positive-vla-bound-value Med
A.5 T2L Toolkit List
We list the tools used in T2L-Agen and their roles in the analysis workflow. The framework is modular and can be easily extended with new tools based on task requirements.
Table 6: T2L Toolkit list and the usage description.
Tool (NAME) Description
view_source Preview a source file with line numbers. Optionally specify start_line/end_line.
grep_source Search code by regex under a root directory. Returns ‘file:line:match‘ lines.
insert_print Insert a single line of debug print before the given line number in a source file.
build_project Build the project inside container. Default workdir=/src.
container_exec Run an arbitrary shell command inside the container (advanced use).
copy_out Copy a file/dir from container to host.
giveup Give up this case to terminate it immediately. Use this to stop solving the ARVO container.
diff_index Parse unified diff and build a simple line-level index without extra anchoring. Output JSON has per-file {anchors_old, insert_points=[], per_line{line->{roles, matched}}}; all line numbers are OLD-file coordinates from the diff.
static_analysis Run comprehensive analysis. For binaries: run Ghidra RE then static tools; for sources: run static tools directly. Uses cppcheck/clang-tidy/infer; aggregates findings to JSON and env state.
chunk_case Parse C/C++ sources under root_dir with tree-sitter and save chunks JSON (index, file_path, chunk_kind, symbol, start/end_line, source, ast_type, imports).
publish_verified_locations Verify/overwrite LLM locations by matching symbol+snippet in numbered snapshots; fallback to original lines; save verified JSON and snapshot index; updates env._state[’last_llm_json_path’].
run_san Run ARVO workflow that triggers ASAN/fuzzer and capture output.
run_gdb Run gdb in the ARVO container and return a backtrace.
llm_analyze Send crash/ASAN log to the LLM to predict likely bug locations (JSON expected). Supports refine mode with source slices.
mark_diff Load anchored diff (anchors_old + insert_points) and mark chunks touched by these lines; updates chunks JSON (adds diff, diff_hit_lines).
compare_llm_metrics Compute detection rate (chunk-level) and localization rate (diff-line-level), plus strict localization by exact interval equality; updates JSON flags accordingly.
gdb_script Run GDB non-interactively with -batch and provided commands; returns GDB output.
extract_json Extract JSON array from a raw LLM response; optionally merge with last predictions; writes temp JSON and updates state.
extract_modified_lines Parse a unified diff and return modified (file, line) list.
compare_patch Compare LLM predicted spans with the patch; set solved by match rate.
pipeline End-to-end: ASAN, GDB, LLM analyze, (optional) compare with patch.
delegate Delegate a task to an executor LLM agent (autonomous, equipped for CTF-style tasks).
static_analysis_config Enable/disable static analysis inclusion in crash log (enabled flag).
baseline_llm_analyze Single-shot baseline without ASAN/GDB or static context—ask LLM to guess once (no refine/verify/postprocess).
A.6 Other case studies
In this section, we present several additional demonstration cases that were not included in the main body of the paper. These examples aim to further illustrate the internal workflow, reasoning strategies, and decision-making processes of the T2L-Agent across diverse scenarios. By showcasing these supplementary cases, we hope to enhance the reader’s understanding of how T2L-Agent performs trace-to-line localization in practice and highlight its robustness across varying bug types and complexity levels.
Figure 13: GPT-4o mini Divergence Tracing for case 12679.
Figure 14: GPT-4o mini Divergence Tracing for case 16857.
Figure 15: GPT-4o mini Divergence Tracing for case 20112.
Figure 16: GPT-4o mini Divergence Tracing for case 20217.
Figure 17: GPT-4o mini Divergence Tracing for case 25357.
Figure 18: GPT-4o mini Baseline for case 12679.